# Quarter- and half-filled quantum Hall states and their competing interactions in bilayer graphene

### v2

@@ Abstract
Figure 4: Quarter-filled states in BLG. (A) Classification of quarter-filled states and their identification based on daughter-state fillings. The pairing channel $\ell$ corresponds to the number of Majorana edge channels (see SM Sec. 13). (B) $R_{xx}$ and $R_{xy}$ around $\nu=\frac{3}{4}$ measured at $D=-76~{}\mathrm{mV/nm}$, along with possible daughter states marked by vertical dashed lines. (C) The thermal activation gap of $\nu=\frac{3}{4}$ measured as a function of $B_{\parallel}$ at $B_{\perp}~{}=~{}15~{}\mathrm{T}$. (D) $\Delta_{\frac{3}{4}}$ measured as a function of $D$ in different valleys. (E) $R_{xx}$ around $\nu=\frac{3}{4}$ measured as a function of $B$ for $D=-70~{}\mathrm{mV/nm}$. (F) Constraints on the ground states of all observed quarter-filled states based on their daughters, along with the measured thermal activation gaps.
@@
To probe the valley polarization, we measured $\Delta_{\frac{3}{4}}$ as a function of $D$; see Fig. 4D and SM Sec. 6 for more details. For the $K^{\prime}$ valley, we find that the gap remains constant as the absolute value of $D$ increases. By contrast, in the $K$ valley, the gap decreases monotonically. However, this behavior does not necessarily imply a valley singlet state. Indeed, a reduction of the gap with $D$ was also reported for $\nu=\frac{3}{2}$.[18] There, the gap eventually saturates to a non-zero value, and a single-component state is further supported by the daughters.[22] Consequently, our measurement can neither rule out an anti-Pfaffian nor a valley-singlet anti-(331) state. Fig. 4E shows $R_{xx}$ as a function of $\nu$ and $B$, with the incompressible state at $\nu=\frac{3}{4}$ persisting down to $B=12~{}\mathrm{T}$. We summarize our analysis of all observed quarter-filled states and their daughters in the table of Fig. 4F (see SM Secs. 5 and 8 for gap measurements and daughters of other quarter-filled states).
Figure 5: Odd- and even-denominator FQH states in partially filled Landau levels of 2CFs. $R_{xx}$ and $R_{xy}$ measured at $B=18~{}\mathrm{T}$ and $D=55~{}\mathrm{mV/nm}$ between $\nu=\frac{1}{3}$ and $\frac{3}{7}$. Dashed vertical and horizontal lines mark the minima in $R_{xx}$ and plateaus in $R_{xy}$, respectively. The electron fillings $\nu=\frac{6}{17},\frac{4}{11},\frac{3}{8}$ correspond to 2CFs at fillings $\nu_{\mathrm{CF}}=\frac{4}{3},\frac{6}{5}$, and $\frac{3}{2}$, respectively. The magnified region in the inset is indicated in Fig. 2A and shows $R_{xx}$ as a function of $\nu$ and $D$. The dotted blue line marks the $R_{xx}$ line cut at $D=55~{}\mathrm{mV/nm}$.
Finally, we have also observed ‘next-generation’ FQH states, which are not captured by weakly interacting CFs.[56, 46, 57, 58, 32] In Fig. 5, we show $R_{xx}$ and $R_{xy}$ at $D=55~{}\mathrm{mV/nm}$ from $\frac{1}{3}\leq\nu\leq\frac{3}{7}$, identifying two odd-denominator states $\nu=\frac{4}{11}$ and $\nu=\frac{6}{17}$ as well as the even denominator $\nu=\frac{3}{8}$ (see SM Sec. 11 for additional next-generation states). They predominantly form in the $N=0$ orbital, similar to the paired 4CF states at $\nu=\frac{3}{4}$. We note that the $\nu=\frac{3}{4}$ states can also be viewed as FQH states of 2CFs at $\nu_{\mathrm{CF}}=-\frac{3}{2}$;[32] see also SM Sec. 13. These observations suggest that the $N=0$ levels of BLG favor the formation of FQH states of 2CFs.
@@
Conclusion. We uncovered a systematic and universal manner in which the topological orders of half-filled FQH states in BLG are determined. Near crossings of two $N=1$ levels, we consistently find (aPf, Pf) in the lower and upper levels, respectively, suggesting that Landau-level mixing within these pairs is decisive. Additionally, we have observed four systematically occurring paired states of 4CFs and constrained their topological order based on daughter states to be of either anti-Pfaffian or anti-(331) order. The quarter-filled states arise exclusively in $N=0$ orbitals, oppositely to the half-filled states, which occur at $N=1$. Moreover, the $N=0$ orbitals also host ‘next-generation’ odd- and even-denominator FQH states, previously only observed in GaAs[56, 46, 58, 59] and in suspended monolayer graphene.[60]
@@
Acknowledgements
It is a pleasure to thank Moty Heiblum and Mike Zaletel for illuminating discussions. Funding: R.K. acknowledges support from the Dean of the Faculty and the Clore Foundation. Y.R. acknowledges the support from the Quantum Science and Technology Program 2021, the Schwartz Reisman Collaborative Science Program, supported by the Gerald Schwartz and Heather Reisman Foundation, the Minerva Foundation with funding from the Federal German Ministry for Education and Research, and the European Research Council Starting Investigator Grant Anyons 101163917. D.F.M. acknowledges support from the Israel Science Foundation (ISF) under grant 2572/21 and from the Minerva Foundation with funding from the Federal German Ministry for Education and Research.
@@ Acknowledgements
See pages 1-23 of Supplementary_material.pdf

### v3

@@ Abstract
Figure 3: Topological orders of half-filled states in BLG. (A) Classification of half-filled states and their identification based on daughter-state fillings. The pairing channel $\ell$ corresponds to the number of Majorana edge channels (see supplementary note 13). (B) $R_{xx}$ and $R_{xy}$ around $\nu=\frac{1}{2}$, measured at $B=18~\mathrm{T}$ and $D=-140~\mathrm{mV/nm}$, along with possible daughter states marked by vertical dashed lines. (C) The thermal activation gap $\Delta_{\frac{1}{2}}$ measured as a function of $B_{\parallel}$ at $B_{\perp}=15~\mathrm{T}$. The dashed line represents the gap value at $B_{\perp}~=~15~\mathrm{T}$ and $B_{\parallel}=0~\mathrm{T}$. (D) Identification of all observed half-filled states based on their daughters, along with the measured thermal activation gaps. The gap of all the half-filled states was measured at $B=18~\mathrm{T}$, except for $\nu=-\frac{3}{2}$ which was measured at $B=11.2~\mathrm{T}$.(E) $R_{xx}$ around $\nu=\frac{3}{2}$ measured as a function of $B$ for $D=-140~\mathrm{mV/nm}$.
Now we turn to a detailed analysis of the quarter-filled states. Similar to their cousins at half filling, quarter states can also exhibit characteristic daughters,[38] yet to be observed in other systems. In Fig. 4A, we show the filling factors of daughters associated with the main pairing channels. In Fig. 4B, we show $R_{xx}$ and $R_{xy}$ around $\nu=\frac{3}{4}$ with the possible daughter states indicated by dashed lines, with their colors referring to the topological order of the parent state from Fig. 4A. The line at $\nu=\frac{19}{25}$ shows remarkable agreement with our data, thereby suggesting the $\nu=\frac{3}{4}$ topological order to be the non-Abelian anti-Pfaffian or the Abelian anti-(331). The absence of the daughter state to the left of $\nu=\frac{3}{4}$ is currently unclear and remains a subject of future investigation.
Figure 4: Quarter-filled states in BLG. (A) Classification of quarter-filled states and their identification based on daughter-state fillings. The pairing channel $\ell$ corresponds to the number of Majorana edge channels (see supplementary note 13). (B) $R_{xx}$ and $R_{xy}$ around $\nu=\frac{3}{4}$ measured at $D=-76~\mathrm{mV/nm}$, along with possible daughter states marked by vertical dashed lines. (C) The thermal activation gap of $\nu=\frac{3}{4}$ measured as a function of $B_{\parallel}$ at $B_{\perp}~=~15~\mathrm{T}$. The dashed line represents the gap value at $B_{\perp}~=~15~\mathrm{T}$ and $B_{\parallel}=0~\mathrm{T}$. (D) $\Delta_{\frac{3}{4}}$ measured as a function of $D$ in different valleys. (E) $R_{xx}$ around $\nu=\frac{3}{4}$ measured as a function of $B$ for $D=-70~\mathrm{mV/nm}$. (F) Constraints on the ground states of all observed quarter-filled states based on their daughters, along with the measured thermal activation gaps.
@@
To determine the spin polarization of the $\nu=\frac{3}{4}$ state, we measured its gap while varying $B_{\parallel}$ at $B_{\perp}=15~\mathrm{T}$; see Fig. 4C and supplementary note 7 for more details. The weak dependence of $\Delta_{\frac{3}{4}}$ on $B_{\parallel}$ up to a Zeeman energy of approximately $14~\mathrm{K}$ again indicates a fully spin-polarized state. In conjunction with the observed daughter state, this observation supports the non-Abelian anti-Pfaffian ground state, as the spin-singlet anti-(331) state should not be spin-polarized.[60] However, a valley-singlet anti-(331) state remains possible. To probe the valley polarization, we measured $\Delta_{\frac{3}{4}}$ as a function of $D$; see Fig. 4D and supplementary note 6 for more details. For the $K^{\prime}$ valley, we find that the gap remains constant as the absolute value of $D$ increases. By contrast, in the $K$ valley, the gap decreases monotonically. However, this behavior does not necessarily imply a valley singlet state. Indeed, a reduction of the gap with $D$ was also reported for $\nu=\frac{3}{2}$.[18] There, the gap eventually saturates to a non-zero value, and a single-component state is further supported by the daughters.[22] Consequently, our measurement can neither rule out an anti-Pfaffian nor a valley-singlet anti-(331) state. Fig. 4E shows $R_{xx}$ as a function of $\nu$ and $B$, with the incompressible state at $\nu=\frac{3}{4}$ persisting down to $B=12~\mathrm{T}$. We summarize our analysis of all observed quarter-filled states and their daughters in the table of Fig. 4F (see supplementary notes 5 and 8 for gap measurements and daughters of other quarter-filled states).
Finally, we have also observed ‘next-generation’ FQH states previously unseen in BLG which are not captured by weakly interacting CFs.[61, 48, 62, 63, 32] Namely, we observed the odd-denominator states $\nu=\frac{4}{11}$ and $\nu=\frac{6}{17}$ and the even-denominator state $\nu=\frac{3}{8}$; see supplementary note 11 for details.
@@
Discussion. We uncovered a systematic and universal manner in which the topological orders of half-filled FQH states in BLG are determined. Near crossings of two $N=1$ levels, we consistently find (aPf, Pf) in the lower and upper levels, respectively, suggesting that LLM within these pairs is decisive. Additionally, we have observed four systematically occurring paired states of 4CFs and suggested their topological order based on daughter states to be of either anti-Pfaffian or anti-(331) order. The quarter-filled states arise exclusively in $N=0$ orbitals, oppositely to the half-filled states, which occur at $N=1$. Moreover, the $N=0$ orbitals also host ‘next-generation’ odd- and even-denominator FQH states, previously only observed in GaAs[61, 48, 63, 64] and in suspended monolayer graphene.[65]
@@
Acknowledgments
It is a pleasure to thank Moty Heiblum and Mike Zaletel for illuminating discussions. Funding: R.K. acknowledges support from the Dean of the Faculty and the Clore Foundation. Y.R. acknowledges the support from the Quantum Science and Technology Program 2021, the Schwartz Reisman Collaborative Science Program, the Gerald Schwartz and Heather Reisman Foundation, a research grant from the Goldfield Family Charitable Trust, the Minerva Foundation with funding from the Federal German Ministry for Education and Research, a research grant from the Estate of Hermine Miller, the Sheba Foundation, and Dweck Philanthropies, Inc., and the funding by the European Union (ERC, Anyons, 101163917). D.F.M. acknowledges support from the Israel Science Foundation (ISF) under grant 2572/21 and from the Minerva Foundation with funding from the Federal German Ministry for Education and Research.
@@ Acknowledgments
See pages 1-30 of Supplementary_material.pdf